Tameen keeps quarters, nickels, and dimes in his change jar. He has three more quarters than dimes and five fewer nickels than dimes. If there are 52 coins in the change jar, how many dimes are in the jar?
Found 2 solutions by stanbon, fractalier:
Answer by stanbon(75887)

Tameen keeps quarters, nickels, and dimes in his change jar. He has three more quarters than dimes and five fewer nickels than dimes. If there are 52 coins in the change jar, how many dimes are in the jar?
Equations:
n + d + q = 52 coins
q = d + 3
n = d-5
----------
Substitute for "n" and "q" and solve for "d"::
d+3 + d + d-5 = 52
3d - 2 = 52
3d = 54
d = 18 (# of dimes)
q = d+3 = 21 (# of quarters)
n = d-5 = 12 (# of nickels)
Cheers,
Stan H.
------------

Answer by fractalier(6550)

Call quarters, nickels, and dimes, q, n and d respectively.
Here is what they tell us...
q + n + d = 52
q = d + 3
n = d - 5
Now substitute these into the first equation and get
(d + 3) + (d - 5) + d = 52
3d - 2 = 52
3d = 54
d = 18 dimes
q = 21 quarters
n = 13 nickels
52 coins total